Ticket Number: HEA-NOC/20051121-2 Ticket Status: CLOSE
Ticket Type: scheduled Resolver: LanComms
Ticket Opened: 20051121 12:25 UTC Problem Start: 20051123 08:00 UTC
Ticket Closed: 20051124 16:36 UTC Problem End: 20051123 10:00 UTC
Site/line: CAO, Limerick PoP
Due to some repeated SPF recalculations Lancomms will be doing some
investigation work on the routers in the Limerick PoP. SPF was being
recalculated ever 10seconds and this is not sustainable.
Some OSPF neighbours will be reset during the investigation which may lead
to short interuptions in inter-PoP traffic and connections on the primary
CAO connection. Traffic will reroute if necessary and any disruption will
20051121 12:20 BN Confirmed affected routers with Lancomms, okayed timing
of work and opened ticket to track.
20051124 16:35 BN This work was completed without service interuption.
Time to Fix:
This ticket can be monitored at http://www.hea.net/tickets/20051121-2